At what distance from the earth, on the line between earth and moon, an object will not feel any gravitational force? Assume earth and moon to be point masses and there is no other force by any other body.                                            Mass of earth-: 6×10^24 kg.                      Mass of moon-: 7.2×10^22kg.                   Average distance between earth and moon is 384000000m.
 

Grade:9

1 Answers

Rama sai rohan
6 years ago
Let us assume that at "x" distance from the earth the object will not feel gravitation . Total distance between moon and earth is 384000000so at a distance "x" Gravitational forces are balanced . Mass of earth is M Mass of moon is MMass of object is s GMs/x^2=Gms/(384000000-x)^2M/x^2= m/(384000000-x)^26×10^24/x^2=7.2×10^22/(384000000-x)^2If you solve this you get x=285202838.6m